OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±8) and asymptotes at y equals inverse of plus minus 1 divided by 2x.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is the center of this hyperbola?

OpenStudy (anonymous):

0,0

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

so (h,k) = (0,0) h = 0 k = 0

OpenStudy (anonymous):

hey that makes a cool face 0,0

OpenStudy (anonymous):

sorry off topic

jimthompson5910 (jim_thompson5910):

lol yes it does how far do go from (0,0) to either vertex?

OpenStudy (anonymous):

8 units

jimthompson5910 (jim_thompson5910):

so a = 8 a = half of the length of the transverse axis the transverse axis connects the two vertices and goes through the center

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

SO what is the formula that we need to plug these into?

jimthompson5910 (jim_thompson5910):

how is the hyperbola oriented? is it opening left/right? or up/down?

OpenStudy (anonymous):

up/down

jimthompson5910 (jim_thompson5910):

according to this http://www.purplemath.com/modules/hyperbola.htm it says that if the hyperbola opens up/down, then the basic form is \[\Large \frac{(y-k)^2}{a^2} + \frac{(x-h)^2}{b^2} = 1\]

OpenStudy (anonymous):

yup

jimthompson5910 (jim_thompson5910):

on that same page it says that the asymptotes have a slope of +-a/b focus on say the positive slope slope = 1/2 a/b = 1/2 8/b = 1/2 b = ??

OpenStudy (anonymous):

16

jimthompson5910 (jim_thompson5910):

\[\Large \frac{(y-k)^2}{a^2} - \frac{(x-h)^2}{b^2} = 1\] \[\Large \frac{(y-0)^2}{8^2} - \frac{(x-0)^2}{16^2} = 1\] \[\Large \frac{y^2}{64} - \frac{x^2}{256} = 1\]

jimthompson5910 (jim_thompson5910):

I had a + in there for some reason. It should have been -

OpenStudy (anonymous):

Ok thanks! I have 4 more questions of similar nature if you dont't mind :)

OpenStudy (anonymous):

I'll make new questions

jimthompson5910 (jim_thompson5910):

try the next one out and I'll take a look at the work you have so far on it

OpenStudy (anonymous):

Not the same type of question but the same chapter in my book